### v1

# Partial fraction decompositions on hyperplane arrangements

Abstract.
We initiate the study of partial fraction decompositions (PFDs) in several variables using tools from commutative algebra. We give criteria for when a rational function with poles on a hyperplane arrangement has a desirable PFD. Our criteria are obtained by examining the primary decomposition of ideals coming from hyperplane arrangements. We then present an algorithm for finding a PFD that satisfies properties desired by physicists, and demonstrate the effectiveness of this algorithm for computing large examples coming from Feynman integrals.
1. Introduction
Partial fraction decompositions (PFDs) of rational functions in several variables appear frequently in physics. Examples include abstract applications, such as the re-casting of scattering amplitudes as canonical forms of polytopes [ABL], and computational applications like developing efficient algorithms to compute Feynman integrals [HvM]. Despite their ubiquity in physics, studies of PFDs seldom appear in mathematical literature [BBD+, Bec, Lei]. There has been interest from physicists to utilize algebro-geometric techniques in order to compute certain PFDs [BW+, CDE, DP]. In this paper we study PFDs from the perspective of commutative algebra, by using primary decomposition and algorithms based upon Gröbner bases.
Let $R=K[x_{1},\ldots,x_{r}]$ be a polynomial ring over a field $K$, and let $f,g\in R$ be nonzero polynomials. We are interested in partial fraction decompositions (PFDs) of the rational function $f/g$; these are expressions of the function as:
$\frac{f}{g}=\sum_{i=1}^{m}\frac{f_{i}}{g_{i}},\qquad\deg(f_{i})<\deg(f),\,\deg(g_{i})<\deg(g).$
Example $1.$
To see where the complexity in PFDs can arise, consider the following example of a rational function with two different PFDs:
$\begin{split}\frac{13y-6x}{(y-3x)(x+y)(x-2y)}&=\frac{19}{12x(x+y)}-\frac{2}{15x(x-2y)}-\frac{33}{20x(y-3x)}\\ &=\frac{1}{(x+y)(x-2y)}+\frac{2}{(y-3x)(x-2y)}-\frac{5}{(y-3x)(x+y)}.\end{split}$
This example demonstrates that:
  1. (1)
    rational functions do not necessarily have a unique PFD; and
  2. (2)
    certain PFDs of rational functions can introduce spurious poles, i.e., poles in the terms of the PFD that do not appear in the rational function itself (see Definition 5.4 for a precise definition).
In applications to physics, one would like to know when a rational function $f/g$ has a PFD, and if so, how to make the PFD “optimal.” Here, optimal can refer to a number of parameters, including minimizing the number of terms $m$ in the PFD, maximizing the degrees of the numerators $f_{i}$, and avoiding spurious poles (the introduction of such poles can complicate calculations and obscure physical properties and symmetries [Hod]).
In this paper, we ask the following question: given a rational function $f/g$ such that $g=\ell_{1}\cdots\ell_{n}$ can be factored into linear polynomials in $R$, can we determine when an optimal PFD for $f/g$ exists? We desire a combinatorial answer in terms of the hyperplane arrangement associated with the linear polynomials $\ell_{1},\ldots,\ell_{n}$. We say that such a rational function $f/g$ has a PFD (of degree $d$) if it can be expressed as:
$\frac{f}{g}=\frac{f}{\ell_{1}\cdots\ell_{n}}=\sum_{\{j_{1},\ldots,j_{n-d}\}\subset[n]}\frac{h_{j_{1}\cdots j_{n-d}}}{\ell_{j_{1}}\cdots\ell_{j_{n-d}}},\qquad\deg(\ell_{i})=1,\,\deg(h_{j_{1}\cdots j_{n-d}})\leq\deg(f)-d.$ (1)
The main results (Theorems 5.2 and 5.3) of this paper characterize the answer to this question in terms of the numerator $f$’s geometric properties relative to the hyperplane arrangement; the main result for hyperplane arrangements in projective space is the following.
Theorem 1.2.
Let $f\in K[x_{1},\ldots,x_{r}]$ be a homogeneous polynomial of degree $\deg(f)\geq d\geq 1$, and let $\ell_{1},\ldots,\ell_{n}$ be nonzero linear forms with $n\geq d$. Then $f/(\ell_{1}\cdots\ell_{n})$ has a PFD of degree $d$ as in (1) if and only if $f$ vanishes to order at least dn+|S|d-n+|S| on all flats $S$ of size at least nd+1n-d+1 of the hyperplane arrangement associated with the linear forms.
Our approach to proving Theorem 1.2 is with commutative algebra: given linear forms $\ell_{1},\ldots,\ell_{n}$ of the polynomial ring $R=K[x_{1},\ldots,x_{r}]$, we study the primary decomposition of the following ideal:
$I_{L,d}=\langle\ell_{i_{1}}\cdots\ell_{i_{d}}:1\leq i_{1}<\cdots<i_{d}\leq n\rangle.$ (2)
Existence of a PFD of order $d$ for the rational function $f/(\ell_{1}\cdots\ell_{n})$ is equivalent to the numerator $f$ being in the ideal $I_{L,d}$, and our primary decomposition gives an equivalent geometric condition for when this occurs.
One can use Macaulay2 [M2] to determine the existence of and algorithmically compute PFDs of the form in (1). Our algorithm (Algorithm 2) for finding a PFD is based upon Gröbner basis algorithms. It satisfies conditions (i), (ii), (iii), and (iv) of the “wishlist” for a good partial fraction algorithm, described by Heller and von Manteuffel in [HvM, §2.3] and based upon work of Leĭnartas [Lei]:
  1. (i)
    the algorithm provides a unique answer;
  2. (ii)
    the algorithm does not introduce spurious denominator factors;
  3. (iii)
    the algorithm commutes with summation;
  4. (iv)
    the algorithm eliminates any spurious poles that are present in the input.
We now discuss the organization and contributions of this paper. In §2, we study the ideals $I_{L,d}$ in (2) defined from hyperplane arrangements. We give primary decompositions for these ideals (Corollary 2.9 and Theorem 2.12), which are presented in terms of the matroid associated with the hyperplane arrangement. Our primary decomposition also allows us to see that the ideals $I_{L,d}$ have linear resolutions (Corollary 2.11). We then apply our decomposition in §3 to study the ideals $I_{L,d}$ where $L$ is the braid arrangement. In §4, we also use our primary decomposition to deduce vanishing behavior of adjoint polynomials arising from polytopes; these polynomials are objects of much study in physics, combinatorics, and algebraic geometry. We then return to our main motivation of PFDs in §5: in this section, we give criteria for when PFDs of degree $d$ exist (Theorems 5.2 and 5.3). We present Algorithms 1 and 2 to determine the existence of and compute PFDs; we also demonstrate how our algorithms satisfy all four conditions (i)(iv) of a good PFD algorithm. In §6, we then show how our criteria and algorithm can be applied to PFDs coming from two different contexts from physics: Feynman integrals (§6.1) and flat-space wavefunction coefficients (§6.2). We conclude in §7 with remarks on the outlook and directions for future investigation for the algebraic study of PFDs.
Code for the algorithms and examples presented in this paper are publicly available at:
2. Primary decompositions
We fix a field $K$, let $r,n\geq 1$ be arbitrary positive integers, and let $R=K[x_{1},\ldots,x_{r}]$ denote the polynomial ring in $r$ variables. The central geometric objects of this paper are hyperplane arrangements, defined by $n$ linear polynomials $L=(\ell_{1},\ldots,\ell_{n})$, $\ell_{i}\in R$; we consider arrangements in both affine and projective spaces, with the $\ell_{i}$ homogeneous in the latter setting.
Let $d$ be a positive integer with $1\leq d\leq n$. Given a hyperplane arrangement $L=(\ell_{1},\ldots,\ell_{n})$, we study the ideal $I_{L,d}\subset R$ defined by $d$-fold products of the linear forms:
$I_{L,d}=\langle\ell_{i_{1}}\cdots\ell_{i_{d}}:1\leq i_{1}<\cdots<i_{d}\leq n\rangle.$
We allow for the cases that some of the $\ell_{i}=0$ are zero, or that the $\ell_{i}$ may not all be distinct, i.e., $\ell_{i}=\ell_{j}$ for for some $i\neq j$. However, we assume that $I_{L,d}\neq 0$ is not the zero ideal, so not all of the $\ell_{i}=0$ are zero. The main results of this section give a primary decomposition for $I_{L,d}$.
2.1. Generic hyperplane arrangements
Before tackling the case of an arbitrary hyperplane arrangement, we focus first on generic arrangements in projective space, i.e., the hyperplanes are in linearly general position; the matroid corresponding to the arrangement is the uniform matroid (see Remark 2.4). Here, we let $n\geq r$ be arbitrary positive integers, and fix d=nr+1d=n-r+1.
To state the following proposition, we introduce some notation. Write each linear form $\ell_{i}\in R$ in the form $\ell_{i}=\sum_{j=1}^{r}a_{ij}x_{j}$, with $a_{ij}\in K$. Define $A=(a_{ij})$ to be the $(n\times r)$-matrix of these coefficients. For $1\leq i_{1}<\cdots<i_{r}\leq n$, denote by $A[i_{1},\ldots,i_{r}]$ the $r$-minor of AA obtained by taking the submatrix of AA with rows labeled by $\ell_{i_{1}},\ldots,\ell_{i_{r}}$. Now, define $\mathcal{M}$ to be the $(\binom{n}{d}\times\binom{n}{d})-$matrix of coefficients, with rows labeled by $d$-fold products $\ell_{i_{1}}\cdots\ell_{i_{d}}$ of the linear forms, and columns labeled by degree $d$ monomials of $R$.
Proposition 2.1.
With the notation set above, we have:
$\det(\mathcal{M})=\pm\prod_{1\leq i_{i}<\cdots<i_{r}\leq n}A[i_{1},\ldots,i_{r}].$ (3)
In particular, if every $r$-subset of the linear forms $\ell_{1},\ldots,\ell_{n}$ is linearly independent, then the set of $d$-fold products of the linear forms is also linearly independent.
Proof.
We may assume that $K$ is algebraically closed. Note that each $A[i_{1},\ldots,i_{r}]$ is irreducible, and so by Hilbert’s Nullstellensatz,
[…]
We now show that $\mathcal{V}(A[i_{1},\ldots,i_{r}])\subset\mathcal{V}(\det(\mathcal{M}))$. Suppose that there exists some relation between the $\ell_{i}$, i.e., for some $\{i_{1},\ldots,i_{r}\}\subset[n]$, there exists $c_{1},\ldots,c_{r}\in K$ not all zero such that:
$\sum_{j=1}^{r}c_{i}\ell_{i_{j}}=0.$ (4)
The existence of such a relation is equivalent to $A[i_{1},\ldots,i_{r}]$ being identically zero. In other words, the tuples $(a_{11},\ldots,a_{nr})\in K^{n\times r}$ for which such a relation (4) exist form $\mathcal{V}(A[i_{1},\ldots,i_{r}])$. Define the set $\{i^{\prime}_{1},\ldots,i^{\prime}_{d-1}\}=[n]\setminus\{i_{1},\ldots,i_{r}\}$. Multiplying each term in (4) by $\ell_{i^{\prime}_{1}}\cdots\ell_{i^{\prime}_{d-1}}$, we find that:
$\sum_{j=1}^{r}c_{i}\ell_{i_{j}}\ell_{i^{\prime}_{1}}\cdots\ell_{i^{\prime}_{d-1}}=0.$ (5)
Each $c_{i}$ in (5) is multiplied by a unique collection of $d$ linear forms. But then we see that the set of $d$-fold products of the $\ell_{i}$ is linearly dependent, and so $\det(\mathcal{M})$ vanishes on $\mathcal{V}(A[i_{1},\ldots,i_{r}])$. Thus $\mathcal{V}(A[i_{1},\ldots,i_{r}])\subset\mathcal{V}(\det(\mathcal{M}))$. This implies that:
$\det(\mathcal{M})=h\cdot\prod_{\begin{subarray}{c}\{i_{1},\ldots,i_{r}\}\subset[n]\\ i_{i}<\cdots<i_{r}\end{subarray}}A[i_{1},\ldots,i_{r}],\qquad h\in K[a_{11},\ldots,a_{nr}].$
Each $A[i_{1},\ldots,i_{r}]$ is a degree $r$ polynomial, and so their product has degree $\binom{n}{r}\cdot r$ in the $a_{ij}$. The entries of the matrix $\mathcal{M}$ are each of degree $d$ in the $a_{ij}$, and so $\det(\mathcal{M})$ has degree $\binom{n}{d}\cdot d$. For n=r+d1n=r+d-1, we have $\binom{n}{r}\cdot r=\binom{n}{d}\cdot d$, so $\det(\mathcal{M})$ and the product of $A[i_{1},\ldots,i_{r}]$ are polynomials of the same degree. This is only possible if $h\in K$. The fact that $h=\pm 1$ is proven in the following lemma (Lemma 2.2). ∎
Lemma $2.$
Following the notation of the proof of Proposition 2.1, we have that $h=\pm 1$:
$\det(\mathcal{M})=\pm\prod_{1\leq i_{i}<\cdots<i_{r}\leq n}A[i_{1},\ldots,i_{r}].$ (6)
Proof.
Consider a diagonal term order on the polynomial ring $K[a_{11},\ldots,a_{nr}]$. Then for any $r$-subset $\{i_{1},\ldots,i_{r}\}\subset[n]$, the leading term (with coefficient) of $A[i_{1},\ldots,i_{r}]$ is
$\operatorname{LT}(A[i_{1},\ldots,i_{r}])=\pm a_{i_{1}1}\cdots a_{i_{r}r}.$
Therefore, the leading term of the RHS of (6) is $\pm m$, where $m$ is the monomial with the exponent of the variable $a_{ij}$ given by:
  • if $i\geq j$: $\binom{i-1}{j-1}\binom{n-i}{r-j}$, which is the number of $r$-sequences $(i_{1},\ldots,i_{r})$ where $1\leq i_{1}<\cdots<i_{r}\leq n$ and $i$ appears in the $j^{\text{th}}$ spot;
  • if $i<j$: 0.
Now consider the matrix $\mathcal{M}$. Its columns are indexed by degree $d$ monomials in $R$, so order its columns by lexicographic order. Its rows are indexed by $r$-subsets of $[n]$, so order these by lexicographic order on $n]^{r}$. Then the leading term $m^{\prime}$ of $\det(\mathcal{M})$ is the product of terms appearing in the diagonal of $\mathcal{M}$. Using the “stars and bars” combinatorial argument, i.e., the correspondence between monomials and $n$-tuples of stars and bars, with $i_{1},\ldots,i_{r}$ corresponding to the indices where stars are placed, one sees that $m^{\prime}$ is the monomial with the exponent of the variable $a_{ij}$ given by:
  • if $i\geq j$: $\binom{i-1}{j-1}\binom{n-i}{r-j}$, which is the number of $n$-tuples of stars and bars where the $i^{\text{th}}$ position is a star appearing after j1j-1 bars and before rjr-j bars;
  • if $i<j$: 0.
Thus, we see that the leading terms of both sides of (6) are the same, up to multiplication by $\pm 1$. This implies the desired result. ∎
Theorem 2.3.
Let $L=(\ell_{1},\ldots,\ell_{n})$ be $n$ linear forms in $R$ such that every $r$-subset of the $\ell_{i}$ is linearly independent, and let d=nr+1d=n-r+1. Then $I_{L,d}=\mathfrak{m}^{d}$, where $\mathfrak{m}=\langle x_{1},\ldots,x_{r}\rangle$ is the homogeneous maximal ideal. In particular, the set $\{\ell_{i_{1}}\cdots\ell_{i_{d}}:1\leq i_{1}<\cdots<i_{d}\leq n\}$ forms a $K$-basis of the vector space of degree $d$ homogeneous polynomials in $R$.
Proof.
By Proposition 2.1, the set $\{\ell_{i_{1}}\ell_{i_{2}}\cdots\ell_{i_{d}}:1\leq i_{1}<\cdots<i_{d}\leq n\}$ is linearly independent, and has cardinality $\binom{n}{d}$. The vector space of homogeneous degree $d$ polynomials in $R$ also has dimension $\binom{n}{d}$, and so the set of $d$-fold products of the $\ell_{i}$ forms a basis for this vector space. Consequently, the ideal $I_{L,d}$ is the same as the ideal generated by all degree $d$ polynomials of the polynomial ring; this latter ideal is exactly $\mathfrak{m}^{d}$. ∎
Remark 2.4.
The description of the linear forms $\ell_{i}$ in Theorem 2.3 can be reformulated from a matroid-theoretic perspective, with the configuration of the hyperplane arrangement corresponding to the uniform matroid of rank $r$ on $n$ elements. Explicitly, setting $\{\ell_{1},\ldots,\ell_{n}\}$ as the ground set, the collection of independent sets is:
$\{\{\ell_{i_{1}},\ldots,\ell_{i_{m}}\}:\{i_{1},\ldots,i_{m}\}\subset[n],\ m\leq r\}.$
The flats of this matroid are the subsets of the ground set of cardinality at most $r$, together with the ground set itself. This foreshadows Corollary 2.9 in the following subsection, where we consider arbitrary hyperplane arrangements.
$2.$ Arbitrary projective case
In this subsection, let $r,n,d$ be positive integers with $n\geq d$. We consider arbitrary hyperplane arrangements determined by $n$ linear forms $L=(\ell_{1},\ldots,\ell_{n})$, with $\ell_{i}\in R=K[x_{1},\ldots,x_{r}]$. Our main result (Corollary 2.9) is a primary decomposition for the ideal $I_{L,d}$.
Our argument follows that of [CH, §3], in which the authors study products of ideals generated by linear forms; we note that although of a similar flavor, our ideal $I_{L,d}$ cannot be written as the product of such ideals.
Definition 2.5.
Let II be a homogeneous ideal in the polynomial ring $R$, and let $\mathfrak{m}=\langle x_{1},\ldots,x_{r}\rangle$ be the homogeneous maximal ideal. We define the saturation $I^{\operatorname{sat}}$ of II to be the ideal:
$I^{\operatorname{sat}}=\{f\in R:f\cdot\mathfrak{m}^{p}\subset I\text{ for some }p\in\mathbb{N}\}.$
The saturation degree $\mathrm{sat}(I)$ of II is then defined to be the smallest integer $q\in\mathbb{N}$ such that for any $f$ homogeneous with $\deg(f)\geq q$:
$f\in I\quad\Leftrightarrow\quad f\in I^{\operatorname{sat}}.$
For $S\subset[n]$, let $I_{S}=\langle\ell_{i}:i\in S\rangle$ be the prime ideal generated by a subset $S$ of the linear forms. Notice that any power of $I_{S}$ is an $I_{S}$-primary ideal; our primary decomposition of $I_{L,d}$ consists of these $I_{S}$-primary ideals.
Lemma 2.6.
We have the following inclusion of ideals:
$I_{L,d}\subset\bigcap_{\begin{subarray}{c}S\subset[n],\\ |S|\geq n-d+1\end{subarray}}(I_{S})^{d-n+|S|}.$
Proof.
Fix $S\subset[n]$ with $|S|\geq n-d+1$. It suffices to show that $I_{L,d}\subset(I_{S})^{d-n+|S|}$; in particular, it is enough to show that each generator $\ell_{i_{1}}\cdots\ell_{i_{d}}$ of $I_{L,d}$ is in $(I_{S})^{d-n+|S|}$. Since $S$ and $\{i_{1},\ldots,i_{d}\}$ are subsets of $[n]$, we have that $|S\cup\{i_{1},\ldots,i_{d}\}|\leq n$. Also, since $|S|\geq n-d+1$, we have that $n\leq d+|S|-1<d+|S|$. Applying these two facts in succession, we find that:
$|S\cap\{i_{1},\ldots,i_{d}\}|=|S|+d-|S\cup\{i_{1},\ldots,i_{d}\}|\geq|S|+d-n>0.$
But then $S\cap\{i_{1},\ldots,i_{d}\}$ is non-empty and has cardinality at least dn+|S|d-n+|S|. Choose any subset of this intersection that has cardinality dn+|S|d-n+|S|, say $\{j_{1},\ldots,j_{d-n+|S|}\}$. Then each $\ell_{j_{m}}$ is a generator in $I_{S}$ and so the products $\ell_{j_{1}}\cdots\ell_{j_{d-n+|S|}}$ along with $\ell_{i_{1}}\cdots\ell_{i_{d}}$ belong in $(I_{S})^{d-n+|S|}$. ∎
Lemma 2.7.
Let II be the following ideal:
$I=\bigcap_{\begin{subarray}{c}S\subset[n],\\ |S|\geq n-d+1\end{subarray}}(I_{S})^{d-n+|S|}.$
Then $\operatorname{sat}(I)\leq d$.
Proof.
We may assume that $I_{[n]}=\mathfrak{m}$; otherwise, one may work in a polynomial ring with fewer variables. Each term appearing in the decomposition of II is primary. In particular, $(I_{S})^{d-n+|S|}$ is $I_{S}$-primary. Then, a primary decomposition for the saturation $I^{\operatorname{sat}}$ is exactly the same as that of II, except for the removal of ideals that are $\mathfrak{m}$-primary. We have that $\mathfrak{m}^{d}$ appears in the decomposition, since $I_{[n]}=\mathfrak{m}$. Furthermore, the exponent dn+|S|=dn+n=$d$-n+|S|=d-n+n=d is the largest exponent appearing in the decomposition of II. Thus, we see that:
$I=\bigcap_{\begin{subarray}{c}S\subset[n],\\ |S|\geq n-d+1\end{subarray}}(I_{S})^{d-n+|S|}=\mathfrak{m}^{d}\cap\bigcap_{\begin{subarray}{c}S\subset[n],\\ |S|\geq n-d+1,\\ I_{S}\neq\mathfrak{m}\end{subarray}}(I_{S})^{d-n+|S|}=\mathfrak{m}^{d}\cap I^{\operatorname{sat}}.$
This implies that II and $I^{\operatorname{sat}}$ agree in degrees at least $d$, and so $\operatorname{sat}(I)\leq d$. ∎
Theorem 2.8.
We have the following equality of ideals:
$I_{L,d}=\bigcap_{\begin{subarray}{c}S\subset[n],\\ |S|\geq n-d+1\end{subarray}}(I_{S})^{d-n+|S|}.$
Proof.
We proceed by induction on both $d$ and $\dim(R)=r$. We may assume that $K$ is algebraically closed and therefore infinite, as well as that $r$ is minimal so that $I_{[n]}=\mathfrak{m}$.
We first consider the base cases for the induction. Assume that $d=1$. Then $I_{L,d}=\langle\ell_{i}:i\in[n]\rangle=I_{[n]}=\mathfrak{m}$ and:
$\bigcap_{\begin{subarray}{c}S\subset[n],\\ |S|\geq n\end{subarray}}(I_{S})^{1-n+|S|}=I_{[n]}=I_{L,d}.$
Now suppose $d$ is arbitrary and $r=1$. Then $\ell_{i}=a_{i}x_{1}$ with $a_{i}\in K$ not all zero. We then see that $I_{L,d}=\langle x_{1}\rangle^{d}=I^{d}_{[n]}$, and so the claim follows.
We now consider $d,r\geq 2$. In Lemma 2.6, we have shown one of the inclusions, and so the remainder of this proof is dedicated to showing the reverse inclusion. For each $j\in[n]$, define the ideal $I_{j}$ by:
[…]
The induction assumption on $d$ allows us to write the following decomposition for each $I_{j}$:
$I_{j}=\bigcap_{\begin{subarray}{c}S\subset[n]\backslash\{j\},\\ |S|\geq n-d+1\end{subarray}}(I_{S})^{d-n+|S|}.$
Then:
$\bigcap_{\begin{subarray}{c}|S|\subset[n],\\ |S|\geq n-d+1\end{subarray}}(I_{S})^{d-n+|S|}=\mathfrak{m}^{d}\cap\bigcap_{j=1}^{n}I_{j}.$
To prove the reverse inclusion, it therefore suffices to show that if $f\in\bigcap_{j=1}^{n}I_{j}$ is homogeneous of degree at least $d$, then $f\in I_{L,d}$.
Since $K$ is infinite, there exists a linear form $y\in R$ such that
  • $y$ is a nonzerodivisor on each $I_{j}^{\operatorname{sat}}$ for which $I_{j}$ is positive-dimensional; and
  • $y$ is not in the $K$-span of the linear forms $\ell_{j}$ for all $j$.
Consider the ideals $(I_{L,d}+\langle y\rangle)/\langle y\rangle$ and $(I_{j}+\langle y\rangle)/\langle y\rangle$ in the quotient ring $R/\langle y\rangle$, which has dimension r1r-1. These ideals are generated by $d$ and (d1)(d-1)-fold products of the images of the $\ell_{i}$. One can then invoke the induction hypothesis on $\dim R$ to see that the image of $f$ in $R/\langle y\rangle$ is in $(I_{L,d}+\langle y\rangle)/\langle y\rangle$. Therefore, $f$ can be written as $f=h+yg$, where $h\in I_{L,d}$ and $g\in R$ is of degree at least d1d-1, and so we may assume that $f=yg$. Then, $f=yg\in I_{j}\subset I_{j}^{\operatorname{sat}}$ for each $j$. Since $y$ is a nonzerodivisor on $I_{j}^{\operatorname{sat}}$, we have that $g\in I_{j}^{\operatorname{sat}}$ as well. The preceding argument holds for all $j$ with $I_{j}$ positive-dimensional; if $I_{j}$ has dimension zero, then $I_{j}^{\operatorname{sat}}=R$, and so $g\in I_{j}^{\operatorname{sat}}$ for all $j$. Applying the induction hypothesis along with Lemma 2.7 (so $\operatorname{sat}(I_{j})\leq d-1$), we have that $g\in I_{j}$. Write $y\in\mathfrak{m}=I_{[n]}$ as $y=\sum_{j=1}^{n}y_{j}$, where $y_{j}\in\langle\ell_{j}\rangle$. Then $f=yg=\sum_{j=1}^{n}y_{j}g\in I_{L,d}$, as $y_{j}g\in\langle\ell_{j}\rangle\cdot I_{j}\subset I_{L,d}$. This completes the proof. ∎
We are now able to give a combinatorial description for a primary decomposition of $I_{L,d}$ in terms of the matroid associated with the subspace arrangement, which we now define. The matroid $M$ is on the ground set $n]=\{1,\ldots,n\}$. The independence sets of this matroid are given by subsets $\{i_{1},\ldots,i_{m}\}\subset[n]$ such that the linear forms $\ell_{i_{1}},\ldots,\ell_{i_{m}}$ are linearly independent in the vector space of linear forms in $R$.
Given such a matroid $M$, we have an associated rank function $\mathrm{rk}(-)$ on $M$: for $S\subset[n]$, we define $\mathrm{rk}(S)$ to be the rank of the $K$-vector space spanned by $\{\ell_{i}:i\in S\}$. For a subset $S\subset[n]$, we define the closure $\mathrm{cl}(S)$ of $S$ to be:
$\mathrm{cl}(S)=\{i\in[n]:\mathrm{rk}(S)=\mathrm{rk}(S\cup\{i\})\}.$
A flat of the matroid is defined to be a subset $S\subset[n]$ such that $S=\mathrm{cl}(S)$.
Corollary 2.9.
Let $M$ denote the matroid corresponding to the linear forms $\ell_{1},\ldots,\ell_{n}$. Then we have the following primary decomposition of the ideal $I_{L,d}$:
$I_{L,d}=\bigcap_{\begin{subarray}{c}S\subset[n],\\ |S|\geq n-d+1,\\ \text{$S$ a flat of $M$}\end{subarray}}(I_{S})^{d-n+|S|}.$
Proof.
Let $S\subset[n]$ with $|S|\geq n-d+1$. Then $I_{S}=I_{\mathrm{cl}(S)}$, and $|\mathrm{cl}(S)|\geq|S|$, so:
$(I_{S})^{d-n+|S|}\supset(I_{\mathrm{cl}(S)})^{d-n+|\mathrm{cl}(S)|}.$
Therefore, it is enough to only take the closed sets in the intersection of Theorem 2.8; the closed sets are exactly the flats of $M$. ∎
We note that the primary decomposition in Corollary 2.9 is not necessarily minimal, as the following example illustrates. One may try to use more refined data from the associated matroid to determine the associated primes and minimal primary decomposition of $I_{L,d}$, such as in [CT, Theorem 3.2].
Example 2.10.
Let $r=3,d=4,n=6$. Consider the matrix of coefficients for the linear forms given by:
$\begin{pmatrix}0&1&1&1&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{pmatrix}$.
Then a minimal primary decomposition of the corresponding ideal $I_{L,d}$ is:
$I_{L,d}=\langle x\rangle^{2}\cap\langle y,z\rangle\cap\langle x,y\rangle^{3}\cap\langle x,z\rangle^{3}.$
Since $S=[n]$ is always a flat, the ideal $\mathfrak{m}^{d}=\langle x,y,z\rangle^{4}$ appears in the primary decomposition given in Corollary 2.9. However, in this example it is redundant and does not appear in the minimal primary decomposition.
Consider now the following slightly modified matrix of coefficients:
$\begin{pmatrix}1&1&1&0&0&0\\ 0&0&0&1&1&0\\ 0&0&0&0&0&1\end{pmatrix}$.
Then the minimal primary decomposition of the corresponding ideal $I_{L,d}$ is:
$I_{L,d}=\langle x\rangle\cap\langle x,y\rangle^{3}\cap\langle x,z\rangle^{2}\cap\langle y,z\rangle\cap\langle x,y,z\rangle^{4},$
which agrees with the primary decomposition given in Corollary 2.9.
We end this subsection with the following result on resolutions and Castelnuovo–Mumford regularity of the ideals $I_{L,d}$. The proof follows the one for [CH, Theorem 3.1] using Lemma 2.7 and Theorem 2.8.
Corollary 2.11.
The ideal $I_{L,d}$ has a linear resolution. In particular, the Castelnuovo–Mumford regularity of $I_{L,d}$ is $d$.
2.3. Affine case
From our work in the homogeneous case, we may now deduce results for primary decompositions of hyperplane arrangements in affine space.
Consider linear polynomials $\ell_{1},\ldots,\ell_{n}\in R$. By homogenizing, we get linear forms $\ell_{1}^{\prime},\ldots,\ell_{n}^{\prime}\in S=K[x_{1},\ldots,x_{r+1}]$. Let $I_{L,d}\subset R$ denote the ideal of $d$-fold products of the $\ell_{i}$, and let $J_{L,d}\subset S$ denote the ideal of $d$-fold products of the $\ell_{i}^{\prime}$. Then:
$I_{L,d}=(J_{L,d}+\langle x_{r+1}-1\rangle)\cap R.$
Therefore, we use Corollary 2.9 to obtain the following primary decomposition of $I_{L,d}$ in $R$:
Theorem 2.12.
Let $\ell_{1},\ldots,\ell_{n}\in R$ be linear polynomials. Then we have the following primary decomposition of $I_{L,d}$:
$I_{L,d}=\bigcap(I_{S})^{d-n+|S|},$
where the intersection is over subsets $S\subset[n]$ such that:
  1. (1)
    $S$ indexes a flat of the hyperplane arrangement associated with the homogenizations of the $\ell_{i}$;
  2. (2)
    the hyperplane at infinity is not in the span of this associated flat, i.e., $x_{r+1}\notin I_{S}$; and
  3. (3)
    $|S|\geq n-d+1$.
Proof.
We retain the notation introduced in the paragraph preceding the statement of the theorem. By Corollary 2.9, we have that:
$I_{L,d}$ $=(J_{L,d}+\langle x_{r+1}-1\rangle)\cap R$
$=\left(\left(\bigcap_{\begin{subarray}{c}\text{$S$ a flat for $\ell^{\prime}_{i}$},\\ |S|\geq n-d+1\end{subarray}}(J_{S})^{d-n+|S|}\right)+\langle x_{r+1}-1\rangle\right)\cap R$
$=\bigcap_{\begin{subarray}{c}\text{$S$ a flat for $\ell_{i}^{\prime}$},\\ |S|\geq n-d+1\end{subarray}}\left((J_{S})^{d-n+|S|}+\langle x_{r+1}-1\rangle\right)\cap R.$
If $J_{S}$ is supported on the hyperplane at infinity, i.e., $x_{r+1}\in J_{S}$, then:
$\left((J_{S})^{d-n+|S|}+\langle x_{r+1}-1\rangle\right)\cap R=R,$
and so the term is redundant in the intersection. Otherwise:
$\left((J_{S})^{d-n+|S|}+\langle x_{r+1}-1\rangle\right)\cap R=(I_{S})^{d-n+|S|},$
which gives the desired equality. ∎
Example 2.13.
Consider the lines defined by:
$\ell_{1}=x,\quad\ell_{2}=y,\quad\ell_{3}=x-1,\quad\ell_{4}=y-1,\quad\ell_{5}=x-y$
in the polynomial ring $K[x,y]$. Their homogenizations in $K[x,y,z]$ are:
$\ell_{1}^{\prime}=\ell_{1},\quad\ell_{2}^{\prime}=\ell_{2},\quad\ell_{3}^{\prime}=x-z,\quad\ell_{4}^{\prime}=y-z,\quad\ell_{5}^{\prime}=x-y.$
A primary decomposition for the ideal $I_{L,d}$ of $(d=3)$-fold products of the $\ell_{i}$ is given by:
$I_{L,d}=\langle x,y\rangle\cap\langle x-1,y-1\rangle.$
This can be obtained from a primary decomposition for the ideal $J_{L,d}$ of 33-fold products of the $\ell_{i}^{\prime}$, removing ideals supported on the hyperplane $z=0$, and then setting $z=1$:
$J_{L,d}=\langle x,y\rangle\cap\langle x-z,y-z\rangle\cap\langle x,y,z\rangle^{3}.$
3. The braid arrangement
We demonstrate our primary decomposition (Corollary 2.9) on braid arrangements, which form an important family of hyperplane arrangements arising across algebra, geometry, topology, and combinatorics (see, e.g., [Fei]). Braid arrangements also appear in physics via denominators of broken Parke–Taylor functions [Fro] and Parke–Taylor forms [BEPV].
Definition 3.1.
The braid arrangement $\mathcal{B}_{r}$ consists of the following set of $\binom{r}{2}$ hyperplanes in $R=K[x_{1},\ldots,x_{r}]$:
$\mathcal{B}_{r}=\{\mathcal{V}(x_{i}-x_{j}):\ i,j\in[r],\ i<j\}.$
The flats of the braid arrangement $\mathcal{B}_{r}$ are in bijection with partitions of the set $[r]$; if $B_{1}\sqcup\cdots\sqcup B_{m}=[r]$, then a flat of the arrangement is defined by the equations:
[…]
For each flat of $\mathcal{B}_{r}$, one may therefore associate a partition $\lambda\vdash r$, i.e., a tuple $\lambda=(\lambda_{1},\ldots,\lambda_{r})$ where $\lambda_{1}+\cdots+\lambda_{r}=r$ and $\lambda_{1}\geq\cdots\lambda_{r}\geq 0$. Then, for a flat associated with the partition $\lambda$, the number of hyperplanes contained in the flat is given by:
$\binom{\lambda_{1}}{2}+\binom{\lambda_{2}}{2}+\cdots+\binom{\lambda_{r}}{2}$.
Table 1 summarizes the flats and their associated partitions for $\mathcal{B}_{5}$.
Partition Number of flats Number of hyperplanes Example flat
$(5)$ $\binom{5}{5}=1$ $\binom{5}{2}=10$ $12,13,14,15,23,24,25,34,35,45$
$(4,1)$ $\binom{5}{4}=5$ $\binom{4}{2}=6$ $12,13,14,23,24,34$
$(3,2)$ $\binom{5}{3}=10$ $\binom{3}{2}+\binom{2}{2}=4$ $12,13,23,45$
$(3,1,1)$ $\binom{5}{3}=10$ $\binom{3}{2}=3$ $12,13,23$
$(2,2,1)$ $\binom{5}{2}+\binom{3}{2}=13$ $\binom{2}{2}+\binom{2}{2}=2$ $12,34$
$(2,1,1,1)$ $\binom{5}{2}=10$ $\binom{2}{2}=1$ $12$
$(1,1,1,1,1)$ 1 0
Table 1. The first column lists all the partitions of 55, the second states the number of flats associated with a given partition, the third gives the number of hyperplanes for each such flat, and the fourth gives an example of such a flat. We use ijij to denote the polynomial $x_{i}-x_{j}$ for brevity.
Example 3.2.
Let $r=5$ and $d=8$, and consider the braid arrangement $\mathcal{B}_{5}$. There are $n=\binom{5}{2}=10$ linear forms in our arrangement $L$. Then Corollary 2.9 gives a primary decomposition of $I_{L,d}$ as an intersection of the following kinds of ideals:
  • the homogeneous maximal ideal $\mathfrak{m}$ to the power of $d=8$, corresponding to the single flat associated with $(5)$;
  • five ideals, each to the power of 44, corresponding to flats associated with $(4,1)$;
  • ten ideals, each to the power of 22, corresponding to flats associated with $(3,2)$;
  • ten ideals, each to the power of 11, corresponding to flats associated with $(3,1,1)$.
One may notice from Table 1 and Example 3.2 that if $\lambda\succeq\mu$ under the dominance order on partitions (recall that $\lambda\succeq\mu$ under the dominance order if and only if $\lambda_{1}+\cdots+\lambda_{i}\geq\mu_{1}+\cdots+\mu_{i}$ for all $i\in[r]$), then the size of a flat associated with $\lambda$ is larger than that of $\mu$. The following result shows that this holds more generally, thus showing that the set of partitions associated with flats appearing in our primary decomposition (Corollary 2.9) is a filter in the poset of partitions.
Proposition $3.$
Suppose $\lambda,\mu\vdash r$ be partitions, with $\lambda\succeq\mu$ under the dominance order. Then $|S_{\lambda}|\geq|S_{\mu}|$, where $S_{\lambda}$ (resp. $S_{\mu}$) denotes any flat associated with $\lambda$ (resp. $\mu$). In particular, if flats corresponding to the partition $\mu$ appear in the primary decomposition of Corollary 2.9, then so do flats corresponding to the partition $\lambda$.
Proof.
Let $\lambda=(\lambda_{1},\ldots,\lambda_{r})$ and $\mu=(\mu_{1},\ldots,\mu_{r})$, with $\lambda_{i},\mu_{i}\geq 0$ for all ii. It suffices to assume that $\lambda\succeq\mu$ is a cover relation. By [Bry, Proposition 2.3], this implies that there exists $i,j\in[r]$ with $i<j$ such that $\lambda_{i}=\mu_{i}+1$, $\lambda_{j}=\mu_{j}-1$, $\lambda_{i}-\lambda_{j}\geq 2$, and $\lambda_{k}=\mu_{k}$ for all $k\neq i,j$. Then $|S_{\lambda}|\geq|S_{\mu}|$ if and only if:
$\binom{\lambda_{i}}{2}+\binom{\lambda_{j}}{2}\geq\binom{\mu_{i}}{2}+\binom{\mu_{j}}{2}=\binom{\lambda_{i}-1}{2}+\binom{\lambda_{j}+1}{2}$.
Consequently:
$\left(\binom{\lambda_{i}}{2}+\binom{\lambda_{j}}{2}\right)-\left(\binom{\lambda_{i}-1}{2}+\binom{\lambda_{j}+1}{2}\right)$ $=\lambda_{i}-\lambda_{j}-1\geq 1,$
and so the desired conclusion holds. ∎
4. Adjoint polynomials
A rapidly developing field at the intersection of physics, combinatorics, and algebraic geometry is the study of positive geometries and canonical forms [Lam, RST]. Central mathematical objects in this study are projective polytopes and their adjoint polynomials. In this section, we highlight how the primary decomposition in Corollary 2.9 can be used to easily deduce vanishing properties of adjoint polynomials. We define this polynomial, and refer the reader to the paper [Gae] for further introduction to the relevant terminology.
Definition 4.1.
Let $P$ be a polytope in $\mathbb{R}^{r-1}$. Let $C\subset\mathbb{R}^{r}$ be the cone over $P$ with rays $V(C)$ corresponding to vertices of $P$. For each simplicial cone $S\in T$ of a triangulation $T$ of $C$, denote by $a_{S}$ the volume of the parallelepiped whose edges are defined by the rays $V(S)$ of $S$. The adjoint polynomial of $C$ is defined as:
$\mathrm{adj}_{C}(x)=\sum_{S\in T}\left(a_{S}\prod_{v\in V(C)\backslash V(S)}\ell_{v}(x)\right).$
Here, $\ell_{v}(x)=v\cdot x$, where $\cdot$ denotes scalar product.
The form of $\mathrm{adj}_{C}$ is independent of the choice of triangulation of $C$, and $\mathrm{adj}_{C}$ is homogeneous and of degree $|V(C)|-\dim(C)$ [Gae, War]. Furthermore, Kohn–Ranestad showed that the residual arrangement of the hyperplane arrangement defining the facets of the dual cone $C^{\vee}$ must vanish on the adjoint polynomial; the adjoint polynomial is also uniquely characterized by certain vanishing conditions on this residual arrangement [BW, KR].
The above definition of adjoint polynomial makes it clear that adjoint polynomials are in ideals of the form $I_{L,d}$, where in this case the hyperplanes defining the ideal are defined by the $\ell_{v}$ from Definition 4.1. Consequently, using the primary decomposition of Corollary 2.9, one can determine certain properties about where adjoint polynomials vanish, without needing to compute any of the $a_{S}$ coefficients or the residual arrangement of the dual cone. Corollary 2.9 implies the following vanishing behavior of adjoint polynomials.
Proposition 4.2.
Let $P\subset\mathbb{R}^{r-1}$ be a polytope, and let $C$ be the cone over $P$. Then the adjoint polynomial $\mathrm{adj}_{C}$ vanishes to order at least dn+|S|d-n+|S| on all flats $S$ of size at least nd+1n-d+1 of the hyperplane arrangement given by the homogenized facet hyperplanes of $P^{\vee}$.
This vanishing condition is different from that of [KR, Theorem 1.5], which guarantees vanishing of $\mathrm{adj}_{C}$ on the residual arrangement of the dual polytope.
[…]
Figure 1. Polytope used to define the adjoint polynomial in Example 4.3.
Example 4.3.
Consider the polytope $P$ shown in Figure 1. Its homogeneous facet hyperplanes can be given by:
$\begin{aligned} \ell_{1}&=x_{2}-x_{4},\\ \ell_{3}&=-2x_{1}+x_{2}+2x_{4},\\ \ell_{5}&=-2x_{1}+x_{2}-2x_{4},\\ \ell_{7}&=x_{2}-2x_{3}+x_{4}.\end{aligned}\quad\quad\begin{aligned} \ell_{2}&=2x_{1}+x_{2}-2x_{4},\\ \ell_{4}&=2x_{1}+x_{2}+2x_{4},\\ \ell_{6}&=x_{3}\end{aligned}$
Each of these can be written in the form $v_{i}\cdot x$ where the $v_{i}$ are the set of vertex rays of a pointed polyhedral cone $C$ (the dual cone associated with $P$). The adjoint polynomial of this cone is thus of the form:
$\text{adj}_{C}(x)=\sum_{1\leq i_{1}<i_{2}<i_{3}\leq 7}a_{i_{1}i_{2}i_{3}}\ell_{i_{1}}\ell_{i_{2}}\ell_{i_{3}},\qquad a_{i_{1}i_{2}i_{3}}\in\mathbb{R}.$
This polynomial is an element of the ideal:
$I_{L,3}=\langle\ell_{i_{1}}\ell_{i_{2}}\ell_{i_{3}}:1\leq i_{1}<i_{2}<i_{3}\leq 7\rangle,$
which, using Corollary 2.9, can be expressed as the primary decomposition:
$I_{L,3}=\bigcap_{\begin{subarray}{c}S\subset[7],\,|S|\geq 5,\\ \text{$S$ a flat of M}\end{subarray}}(I_{S})^{|S|-4}=\langle\ell_{i}:i\in[5]\rangle\cap\langle\ell_{i}:i\in[7]\rangle^{3}.$
Here, we have used that the matroid defined by the facet hyperplanes only has two flats containing more than four elements, those being $\{\ell_{i}:i\in[5]\}$ and $\{\ell_{i}:i\in[7]\}$. This result tells us that the adjoint polynomial associated with the polytope $P$ vanishes with order at least one where the five linear forms $\ell_{1},\ldots,\ell_{5}$ all vanish.
Remark $4.$
It was shown in [Gae] that canonical forms of polytopes $P$ can be expressed as rational functions, the numerator being the adjoint polynomial associated with the polytope $P^{\vee}$ dual to $P$, and the denominator being a product of facet hyperplanes of $P^{\vee}$. This means that canonical forms are exactly a type of rational function with a PFD of degree $d$.
5. Partial fraction decompositions and algorithmic aspects
We now return to the problem of existence and computation of PFDs. Let $f/(\ell_{1}\cdots\ell_{n})$ be a rational function, where $f,\ell_{1},\ldots,\ell_{n}\in R=K[x_{1},\ldots,x_{r}]$ are nonzero polynomials such that $\deg(\ell_{i})=1$. Recall that we are interested in when this function has a PFD of degree $d$, meaning that it can be expressed as:
$\frac{f}{\ell_{1}\cdots\ell_{n}}=\sum_{\{j_{1},\ldots,j_{n-d}\}\subset[n]}\frac{h_{j_{1}\cdots j_{n-d}}}{\ell_{j_{1}}\cdots\ell_{j_{n-d}}},\qquad\deg(h_{j_{1}\cdots j_{n-d}})\leq\deg(f)-d.$ (7)
Generally, the least common multiple of all of the terms in the PFD expression is exactly $\ell_{1}\cdots\ell_{n}$. By putting all of the terms of the PFD under this common denominator, we see that such a PFD exists exactly when the polynomial $f$ can be expressed as an $R$-linear combination of $d$-fold products of the $\ell_{i}$’s. In other words, a PFD of degree $d$ exists for $f/(\ell_{1}\cdots\ell_{n})$ if and only if $f$ lies in the ideal $I_{L,d}$ studied in §2. We can thus interpret our results on primary decompositions of $I_{L,d}$ as vanishing criteria for existence of PFDs. We first consider the case of hyperplane arrangements in linearly general position.
Theorem 5.1.
Let $\ell_{1},\ldots,\ell_{n}$ be $n$ linear forms in $R=K[x_{1},\ldots,x_{r}]$ such that any $r$-subset of the $\ell_{i}$ are linearly independent. Set d=nr+1d=n-r+1, and let $f\in R$ be any homogeneous degree $d$ polynomial. Then the rational function $f/(\ell_{1}\cdots\ell_{n})$ has a unique PFD of degree $d$.
Proof.
By Theorem 2.3, the polynomial $f$ can be written uniquely as a $K$-linear combination of the elements of $\{\ell_{i_{1}}\ell_{i_{2}}\cdots\ell_{i_{d}}:1\leq i_{1}<\cdots<i_{d}\leq n\}$:
$f=\sum_{\begin{subarray}{c}\{i_{1},\ldots,i_{d}\}\subset[n],\\ i_{1}<\cdots<i_{d}\end{subarray}}a_{i_{1}\ldots i_{d}}\ell_{i_{1}}\cdots\ell_{i_{d}},\qquad a_{i_{1}\cdots i_{d}}\in K.$
Thus, we can write the rational function as:
$\frac{f}{\ell_{1}\ell_{2}\cdots\ell_{n}}=\sum_{\begin{subarray}{c}\{i_{1},\ldots,i_{d}\}\subset[n],\\ i_{1}<\cdots<i_{d}\end{subarray}}\frac{a_{i_{1}\cdots i_{d}}}{(\ell_{1}\ell_{2}\cdots\ell_{n})/(\ell_{i_{1}}\cdots\ell_{i_{d}})}=\sum_{\begin{subarray}{c}\{j_{1},\ldots,j_{n-d}\}\subset[n],\\ j_{1}<\cdots<j_{n-d}\end{subarray}}\frac{b_{j_{1}\cdots j_{n-d}}}{\ell_{j_{1}}\cdots\ell_{j_{n-d}}},$
where we set $b_{j_{1}\cdots j_{n-d}}$ to be equal to $a_{i_{1}\cdots i_{d}}$ with $\{i_{1},\ldots,i_{d}\}=[n]\setminus\{j_{1},\ldots,j_{n-d}\}$. ∎
We now turn to the case of arbitrary hyperplane arrangements, both in the projective and affine settings. The following results on PFDs follow from Corollary 2.9 and Theorem 2.12.
Theorem 5.2.
Let $\ell_{1},\ldots,\ell_{n}$ be $n$ nonzero linear forms in $R$, let $d\leq n$, and let $f\in R$ be a homogeneous polynomial with $\deg(f)\geq d$. Then $f/(\ell_{1}\cdots\ell_{n})$ has a PFD of degree $d$ if and only if $f$ vanishes to order at least dn+|S|d-n+|S| on all flats $S$ of size at least nd+1n-d+1 of the hyperplane arrangement associated with the linear forms.
Theorem 5.3.
Let $\ell_{1},\ldots,\ell_{n}$ be $n$ nonzero linear polynomials in $R$, let $d\leq n$, and let $f\in R$ be a polynomial with $\deg(f)\geq d$. Then $f/(\ell_{1}\cdots\ell_{n})$ has a PFD of degree $d$ if and only if $f$ vanishes to order at least dn+|S|d-n+|S| on all flats $S$, where:
  • $S$ is a flat of the hyperplane arrangement associated with the homogenizations of the $\ell_{i}$;
  • the size of $S$ is at least nd+1n-d+1; and
  • $S$ is not supported on the hyperplane at infinity.
We now turn to the matter of computing PFDs. Recall the wishlist for a good PFD algorithm outlined in §1. We first focus on the notion of spurious poles.
Definition 5.4.
A spurious pole of a rational function is a denominator factor that can be removed upon a re-writing of the rational function.
In Example 1.1, the first decomposition of our rational function introduces a spurious pole at $x=0$; this is undesirable. Our notion of a PFD of degree $d$ guarantees that spurious poles are not introduced, as all of the poles in (7) appear in both the original rational function and its PFD. However, before beginning any computation of PFDs of this form, we must ensure that spurious poles do not already exist in the input of $f/(\ell_{1}\cdots\ell_{n})$. To address this, consider the following definition: we say that a rational function $f/g$ with $f,g\in R$ nonzero polynomials is reducible if $f,g$ share a common irreducible factor. Then, the input rational function $f/(\ell_{1}\cdots\ell_{n})$ already having a spurious pole is equivalent to it being reducible. One also sees immediately that this is equivalent to $f\in\langle\ell_{i}\rangle$ for some $i\in[n]$. Algorithm 1 below constructs a reduced expression for the rational function by finding and eliminating these spurious $\ell_{i}$ (if they exist).
Algorithm 1 ReducedExp
Input: rational function $f/(\ell_{1}\cdots\ell_{n})$ where $f,\ell_{1},\ldots,\ell_{n}\in K[x_{1},\ldots,x_{r}]$ nonzero and $\deg(\ell_{i})=1$
Output: fully reduced expression for $f/(\ell_{1}\cdots\ell_{n})$
[…]
2:for $i$ from 11 to $n$ do
3:  if $f\in\langle\ell_{i}\rangle$ then
4:   write $f=g\ell_{i}$ using //
5:   $f\leftarrow g$
6:  else
7:   add $\ell_{i}$ to $L$
8:  end if
9:end for
10:return $f,L$, where the reduced rational function is given by $f/(\ell_{i_{1}}\cdots\ell_{i_{m}})$ with $\ell_{i_{j}}\in L$
Once a reduced expression for $f/(\ell_{1}\cdots\ell_{n})$ is obtained, one can proceed with the main PFD algorithm (Algorithm 2) to determine if the rational function has a PFD of degree $d$, with $d$ as large as possible. If such a PFD exists, then the algorithm returns one.
Algorithm 2 PFD
Input: fully reduced rational function $f/(\ell_{1}\cdots\ell_{n})$ where $f,\ell_{1},\ldots,\ell_{n}\in K[x_{1},\ldots,x_{r}]$ nonzero and $\deg(\ell_{i})=1$
Output: PFD of degree $d\in[n]$ for $f/(\ell_{1}\cdots\ell_{n})$ with $d$ maximum, if one exists
[…]
3:for $D$ from 11 to $n$ do
4:  create ideal $I_{L,D}$
5:  if $f\in I_{L,D}$ then
6:   $d\leftarrow D$
7:   $C\leftarrow$ coefficients for quotient of $f$ with respect to generators of $I_{L,D}$ using // in M2
8:  else
9:   break
10:  end if
11:end for
12:if $d=0$ then
13:  return no PFD of any degree possible
14:else
15:  return PFD of degree $d$ in terms of $C$
16:end if
Remark 5.5 (Wishlist).
We now explain why our algorithms satisfy the wishlist from §1.
  • (i)
    Uniqueness of output: Given a polynomial $f$ and an ordering $\ell_{1},\ldots,\ell_{n}$ of the denominator factors of the rational function, our algorithm is deterministic in the output – there are no probabilistic aspects of the algorithm. If we are in the case of a generic hyperplane arrangement and $f$ is homogeneous of degree $d$ (as in Theorem 5.1), then the output is also independent of any relabeling of the linear forms.
  • (ii)
    Does not introduce spurious poles: This is guaranteed by the form of our output PFD (1), where denominator factors of the PFD are all subsets of the original $\ell_{i}$.
  • (iii)
    Commutes with summation: This is guaranteed by the use of the // command in Macaulay2. This function uses polynomial division based on Gröbner bases, i.e., expressing polynomials in normal form with respect to the reduced Gröbner basis of the ideal $I_{L,d}$. The map sending polynomials to their normal form is a $K$-linear map.
  • (iv)
    Eliminates any spurious poles present in the input: This is guaranteed by Algorithm 1.
Remark 5.6.
The algorithm above is highly customizable. Below we list examples of small changes that can be made to hone the algorithm for certain applications.
  1. (1)
    Large $n$ or $r$: Construction of the ideals $I_{L,d}$ in Macaulay2 can take a long time when $n$ or $r$ is large. Below are two ways these large computation times can be circumvented.
    1. (a)
      For a given choice of $d$, check if the numerator is in an ideal generated by some subset of the generators of $I_{L,d}$. If this is true, a PFD of degree $d$ is then obtainable with this restricted set of generators. This strategy is used below for the calculation of a “large” coefficient coming from Feynman integrals (§6.1.2).
    2. (b)
      Suppose the numerator is in the ideal $I_{L,d}$ for some small choice of $d$. Rather than checking if the same numerator is in $I_{L,d}$ for larger $d$, consider the numerators in the PFD expansion for small $d$ and apply the PFD algorithm on them, also restricting to small choices of $d$ (if need be). Proceed iteratively. This iterative procedure will produce the same output as the original algorithm as a consequence of the fact that the algorithm commutes with summation (iii).
  2. (2)
    Choosing PFD denominators: Suppose one has a specific PFD form in mind, e.g. that does not have certain denominator factors. This can be achieved by restricting the generators of $I_{L,d}$ to the subset of them that do not produce the undesired denominator factors. If the numerator is a member of this restricted ideal, then the desired PFD can be generated.
  3. (3)
    Nonlinear denominator factors: The algorithm above can readily be applied to the case where the $\ell_{i}$ are not linear. Unlike when all the $\ell_{i}$ are linear, the resulting PFD may contain terms with denominators of differing degrees.
6. Towards physics
In this section, we apply our algorithm from §5 to compute PFDs of rational functions arising from two different contexts in physics: Feynman integrals (§6.1) and flat-space wavefunction coefficients (§6.2). The code for these examples can be found here.
6.1. Feynman integrals
In preparation for the approaching upgrade to the Large Hadron Collider [AAA+], much work has been done in developing efficient algorithms to aid the computation of Feynman integrals [HvM, BFP, CGM]. Many of these algorithms are based on integration-by-parts (IBP) reductions, in which a Feymann integral is decomposed as a linear combination of master integrals (MIs). These MIs are a choice of basis for the family of Feynman integrals under consideration. The coefficients of the linear combination are rational functions in the external kinematics and the dimensional regularization parameter $\epsilon$. These rational functions are often very large, and can make calculations cumbersome. The authors of [BW+] suggest storing the coefficients of IBP reductions in the form of a PFD, and introduce their own PFD algorithm as a “powerful method to reduce the byte size of the analytic IBP reduction coefficients”. Their paper includes an application to IBP reduction coefficients arising from the Feynman integral associated with the the two-loop five-point non-planar double pentagon diagram shown in Figure 2. These coefficients were calculated (with respect to a specific choice of MIs) in [BBD+]. The authors of [HvM] suggest using these coefficients as a benchmark for testing PFD algorithms for Feynman integrals. In this section we naively apply our PFD algorithm to two of these coefficients, a “small” and a “large” example, and discuss the results. We hope this may serve as a starting point for further refinement by physicists.
[…]
    $p_{1}$$p_{2}$$p_{4}$$p_{5}$$p_{3}\endlxSVG@picture$
Figure 2. Two-loop five-point non-planar double pentagon Feynman diagram with labelled external momenta $p_{i}$.
6.1.1. Small IBP coefficient
This IBP coefficient is a rational function in six variables. These variables are the dimensional regulatization parameter $\epsilon$ and the symbolic Mandelstam variables $s_{12}$, $s_{15}$, $s_{23}$, $s_{34}$, and $s_{45}$ (Lorentz invariant functions of the external momenta $p_{i}$). Explicitly, the polynomial ring here is $\mathbb{Q}[s_{12},s_{15},s_{23},s_{34},s_{45},\epsilon]$. The numerator is a degree $11$ polynomial, and the denominator can be factorized as a product of the hyperplanes:
$\begin{aligned} \ell_{1}&=8(1+4\epsilon)\\ \ell_{4}&=s_{12}+s_{15}-s_{34}\\ \ell_{7}&=s_{12}+s_{23}-s_{45}\\ \ell_{10}&=s_{45}\\ \end{aligned}\quad\quad\begin{aligned} \ell_{2}&=s_{12}\\ \ell_{5}&=-s_{15}+s_{23}+s_{34}\\ \ell_{8}&=-s_{15}+s_{23}-s_{45}\\ \ell_{11}&=s_{45}\\ \end{aligned}\quad\quad\begin{aligned} \ell_{3}&=s_{23}\\ \ell_{6}&=s_{12}-s_{45}\\ \ell_{9}&=s_{12}-s_{34}-s_{45}\\ \ell_{12}&=s_{34}+s_{45}.\end{aligned}$
We find that the input rational function contains no spurious denominator factors and that the largest $d$ for which the numerator is in the ideal $I_{L,d}$ is 88. This implies the existence of a PFD of degree 88 as in (7), with degree 44 denominators. Algorithm 2 produces such a PFD with $173$ terms (where the number of distinct terms is the the number of distinct denominators appearing in the PFD). These results were computed within seconds.
6.1.2. Large IBP coefficient
This IBP coefficient is a rational function in the same six variables as above. To reduce computation time, one of the Mandelstam variables can be set to 11. This variable can later be re-introduced through dimensional arguments. We choose to set $s_{12}=1$, and so the polynomial ring here is $\mathbb{Q}[s_{15},s_{23},s_{34},s_{45},\epsilon]$. The numerator is then a degree $30$ polynomial in five variables, and the denominator can be written as a product of the following $29$ hyperplanes:
[…]
Given that we have set $s_{12}=1$, we have suppressed the $30$th denominator hyperplane $\ell_{30}=s_{12}$. Without making any modifications to Algorithm 2, the time it takes to construct the ideals $I_{L,d}$ grows exponentially with $d$, already taking hours at $d=7$. By choosing subsets of the generators of the ideals one can dramatically reduce computation time. The only drawback is that the numerator not being in a subset of $I_{L,d}$ does not rule out the non-existence of a PFD of degree $d$. Using this strategy, we have found that the numerator is in an ideal contained in the ideal $I_{L,17}$. This smaller ideal was generated using $5985$ of the $51895935$ generators of $I_{L,17}$ (we did not include $\ell_{22},\ldots,\ell_{29}$) and took $194$ seconds to construct. Testing that the numerator is in this ideal took $930$ seconds, and generating the corresponding PFD took $2217$ seconds. This PFD has $546$ terms (counted by distinct denominators) and degree $12$ denominators. Comparing with the times given in [HvM, Table 2], our algorithm implemented in Macaulay2 is competitive with other modern PFD algorithms for Feynman integrals.
6.2. Wavefunction coefficients
As mentioned in §1, one of the ambitions of modern theoretical physics is to develop an alternative picture of scattering amplitudes, one that moves away from Feynman integrals and takes a more geometric approach. Landmark papers in this direction include [AT], and show that certain scattering amplitudes can be represented in terms of canonical forms of polytopes such as the amplituhedron and associahedron. Similar developments have been made in the study of cosmological correlators, relating flat-space wavefunction coefficients and canonical forms of cosmological polytopes [ABP]. Below we show an example of how Algorithm 2 can be used to shed light on the structure of flat-space wavefunction coefficients.
[…]
       $Y_{34}$$Y_{24}$$Y_{14}$$X_{4}$$X_{3}$$X_{2}$$X_{1}\endlxSVG@picture$
Figure 3. Graph corresponding to the flat-space wavefunction coefficient $\psi$.
Example 6.1.
As explained in [ABH+], the flat-space wavefunction coefficient $\psi$ associated with the graph shown in Figure 3 can be generated by considering tubings of the graph. The result is the following expression for $\psi$:
$\psi=\frac{8Y_{14}Y_{24}Y_{34}}{\ell_{1}\ell_{2}\ell_{3}\ell_{4}\ell_{5}}\bigg[\frac{1}{\ell_{9}}\bigg(\frac{1}{\ell_{6}}+\frac{1}{\ell_{7}}\bigg)+\frac{1}{\ell_{10}}\bigg(\frac{1}{\ell_{6}}+\frac{1}{\ell_{8}}\bigg)+\frac{1}{\ell_{11}}\bigg(\frac{1}{\ell_{7}}+\frac{1}{\ell_{8}}\bigg)\bigg],$
$\begin{aligned} \ell_{1}&=X_{1}+Y_{14}\\ \ell_{3}&=X_{3}+Y_{34}\\ \ell_{5}&=X_{1}+X_{2}+X_{3}+X_{4}\\ \ell_{7}&=X_{2}+X_{4}+Y_{14}+Y_{34}\\ \ell_{9}&=X_{1}+X_{2}+X_{4}+Y_{34}\\ \ell_{11}&=X_{2}+X_{3}+X_{4}+Y_{14}.\\ \end{aligned}\quad\quad\begin{aligned} \ell_{2}&=X_{2}+Y_{24}\\ \ell_{4}&=X_{4}+Y_{14}+Y_{24}+Y_{34}\\ \ell_{6}&=X_{1}+X_{4}+Y_{24}+Y_{34}\\ \ell_{8}&=X_{3}+X_{4}+Y_{14}+Y_{24}\\ \ell_{10}&=X_{1}+X_{3}+X_{4}+Y_{24}\end{aligned}$
The function $\psi/(8Y_{14}Y_{24}Y_{34})$ is the canonical form of the cosmological polytope whose facet hyperplanes are the denominator factors. In [FPSW], the authors conjecture a formula for the PFD of flat-space wavefunction coefficients associated with arbitrary tree-level cosmological Feynman diagrams. An amended version of this conjecture was recently proven in [Dun]. That a PFD of this form exists in this case is clear from the fact that the numerator of $\psi$ (once all the terms have been collected under a common denominator) is in the ideal:
$I_{L,7}=\langle\ell_{i_{1}}\cdots\ell_{i_{7}}:1\leq i_{1}<i_{2}<\cdots<i_{7}\leq 11\rangle,$
(as can be checked using Macaulay2), and that the numerator is homogeneous of degree 77. And so, using the command // one can express $\psi$ in terms of the generators of this ideal, yielding:
$\begin{split}\psi&=\frac{1}{\ell_{7}\ell_{9}\ell_{10}\ell_{11}}+\frac{1}{\ell_{5}\ell_{9}\ell_{10}\ell_{11}}+\frac{1}{\ell_{4}\ell_{8}\ell_{10}\ell_{11}}+\frac{1}{\ell_{4}\ell_{7}\ell_{10}\ell_{11}}+\frac{1}{\ell_{4}\ell_{7}\ell_{9}\ell_{10}}+\frac{1}{\ell_{4}\ell_{6}\ell_{9}\ell_{10}}+\frac{1}{\ell_{3}\ell_{8}\ell_{10}\ell_{11}}\\ &\ \ \ +\frac{1}{\ell_{3}\ell_{5}\ell_{10}\ell_{11}}-\frac{1}{\ell_{3}\ell_{4}\ell_{7}\ell_{9}}-\frac{1}{\ell_{3}\ell_{4}\ell_{6}\ell_{9}}+\frac{1}{\ell_{2}\ell_{7}\ell_{9}\ell_{11}}+\frac{1}{\ell_{2}\ell_{5}\ell_{9}\ell_{11}}-\frac{1}{\ell_{2}\ell_{4}\ell_{8}\ell_{10}}-\frac{1}{\ell_{2}\ell_{4}\ell_{6}\ell_{10}}\\ &\ \ \ -\frac{1}{\ell_{2}\ell_{3}\ell_{8}\ell_{10}}-\frac{1}{\ell_{2}\ell_{3}\ell_{7}\ell_{9}}+\frac{1}{\ell_{2}\ell_{3}\ell_{5}\ell_{11}}+\frac{1}{\ell_{2}\ell_{3}\ell_{4}\ell_{6}}+\frac{1}{\ell_{1}\ell_{6}\ell_{9}\ell_{10}}+\frac{1}{\ell_{1}\ell_{5}\ell_{9}\ell_{10}}-\frac{1}{\ell_{1}\ell_{4}\ell_{8}\ell_{11}}\\ &\ \ \ -\frac{1}{\ell_{1}\ell_{4}\ell_{7}\ell_{11}}-\frac{1}{\ell_{1}\ell_{3}\ell_{8}\ell_{11}}-\frac{1}{\ell_{1}\ell_{3}\ell_{6}\ell_{9}}+\frac{1}{\ell_{1}\ell_{3}\ell_{5}\ell_{10}}+\frac{1}{\ell_{1}\ell_{3}\ell_{4}\ell_{7}}-\frac{1}{\ell_{1}\ell_{2}\ell_{7}\ell_{11}}-\frac{1}{\ell_{1}\ell_{2}\ell_{6}\ell_{10}}\\ &\ \ \ +\frac{1}{\ell_{1}\ell_{2}\ell_{5}\ell_{9}}+\frac{1}{\ell_{1}\ell_{2}\ell_{4}\ell_{8}}-\frac{1}{\ell_{1}\ell_{2}\ell_{3}\ell_{11}}-\frac{1}{\ell_{1}\ell_{2}\ell_{3}\ell_{10}}-\frac{1}{\ell_{1}\ell_{2}\ell_{3}\ell_{9}}+\frac{1}{\ell_{1}\ell_{2}\ell_{3}\ell_{8}}+\frac{1}{\ell_{1}\ell_{2}\ell_{3}\ell_{7}}\\ &\ \ \ +\frac{1}{\ell_{1}\ell_{2}\ell_{3}\ell_{6}}+\frac{1}{\ell_{1}\ell_{2}\ell_{3}\ell_{5}}-\frac{1}{\ell_{1}\ell_{2}\ell_{3}\ell_{4}}.\end{split}$
Note that this is not the PFD one arrives at for the diagram 2 using the process outlined in [Dun], though both PFDs have the same number of terms. This means that there can be multiple different PFDs of tree-level flat space wavefunction coefficients which have the same number of terms, and in which all numerators are $\pm 1$ and all denominators have the same degree. In other words, such “minimal” flat-space wavefunction coefficient PFDs are not unique.
One may also wonder about the vanishing properties of the numerator. Using the Matroids package for Macaulay2, one finds that the matroid defined by lines $\ell_{i}$ has $88$ flats of five elements, $44$ of six elements, $12$ of seven elements, 11 of eight elements, 33 of nine elements, and 33 of ten elements. Corollary 2.9 then tells us that the numerator vanishes on each of these flats $S$, with order |S|4|S|-4. For example, the flats of ten elements are:
$\begin{split}&\{\ell_{2},\ell_{3},\ell_{4},\ell_{5},\ell_{6},\ell_{7},\ell_{8},\ell_{9},\ell_{10},\ell_{11}\},\\ &\{\ell_{1},\ell_{3},\ell_{4},\ell_{5},\ell_{6},\ell_{7},\ell_{8},\ell_{9},\ell_{10},\ell_{11}\},\\ &\{\ell_{1},\ell_{2},\ell_{4},\ell_{5},\ell_{6},\ell_{7},\ell_{8},\ell_{9},\ell_{10},\ell_{11}\}.\end{split}$
7. Future directions and outlook
We present some potential directions for future work regarding PFDs.
In the generic case, PFDs of degree $d$ are unique (Theorem 5.1). But given arbitrary $f\in R$ and an arbitrary hyperplane arrangement $L$, there can be be many different such PFDs. This is because the products of $d$ of the $\ell_{i}$ need not be algebraically independent. To see this explicitly, consider the following example.
Example 7.1.
Define the hyperplanes $\ell_{1}=x-y,\ \ell_{2}=y+z,\ \ell_{3}=z,\ \ell_{4}=x+y\in K[x,y,z]$, and the polynomial $f$:
$f$ $=2x^{2}y^{2}-2y^{4}+12x^{2}yz+4xy^{2}z-2y^{3}z+10x^{2}z^{2}+4xyz^{2}$
$=3x\ell_{1}\ell_{2}\ell_{3}+2y\ell_{1}\ell_{2}\ell_{4}+7x\ell_{2}\ell_{3}\ell_{4}$
$=(-2x-5y)\ell_{1}\ell_{2}\ell_{3}+2y\ell_{1}\ell_{2}\ell_{4}+10\ell_{1}\ell_{2}\ell_{3}\ell_{4}+(2x+5y)\ell_{2}\ell_{3}\ell_{4}.$
We then have the following two PFDs of degree 33:
$\frac{f}{\ell_{1}\ell_{2}\ell_{3}\ell_{4}}=\frac{3x}{x+y}+\frac{2y}{z}+\frac{7x}{x-y}=-\frac{2x+5y}{x+y}+\frac{2y}{z}+10+\frac{2x+5y}{x-y}.$ (8)
Our algorithm gives a unique partial fraction decomposition, but changing the choice of labeling or ordering of the $\ell_{i}$ could result in different PFDs.
Question 7.2.
Given a hyperplane arrangement $L$ and a polynomial $f\in I_{L,d}$, can one determine the number of unique PFDs of degree $d$?
Another question one can ask concerns when a term in a PFD is reducible, such as in the second PFD of (8), where the term $10$ appears.
Question 7.3.
What are necessary or sufficient conditions for when a term of a PFD of degree $d$ is reducible, or even a polynomial?
A first step in answering this question is the following result.
Lemma 7.4.
Suppose $d=n-1\geq 1$, all of the $\ell_{i}$ are pairwise prime linear polynomials, and $f\in I_{L,d}$ is a polynomial. Then a PFD of degree $d$ for $f/(\ell_{1}\cdots\ell_{n})$ has a reducible term if and only if $f\in\langle\ell_{i}\rangle$ for some $i\in[n]$.
Proof.
Since $f\in I_{L,d}$, it can be written in the form $f=\sum_{j=1}^{n}a_{j}\ell_{1}\cdots\hat{\ell}_{j}\cdots\ell_{n}$ for some $a_{j}\in R$. Then we have the following chain of equivalences:
$\ \ \ \ \ f\in\langle\ell_{i}\rangle\text{ for some }i\in[n]$
$\Leftrightarrow f=\ell_{i}g\text{ for some }g\in R$
$\Leftrightarrow\sum_{j=1}^{n}a_{j}\ell_{1}\cdots\hat{\ell}_{j}\cdots\ell_{n}=\ell_{i}g$
[…]
$\Leftrightarrow\frac{f}{\ell_{1}\cdots\ell_{n}}=h+\sum_{j=1,j\neq i}^{n}\frac{a_{j}}{\ell_{j}}.$
Through our study of ideals of the form $I_{L,d}$, we have proven a number of results about PFDs of rational functions on hyperplane arrangements. This led to new ways of understanding the geometry of adjoint polynomials and canonical forms, as well as a novel algorithmic approach for constructing PFDs of a tailored form. We have demonstrated how our findings have direct applications to a variety of problems in physics and hope that this paper inspires further study of PFDs by algebraists.
Acknowledgements
We thank Bernd Sturmfels, Andreas von Manteuffel, and Julian Weigert for valuable feedback and guidance during the development of this project. CdK was supported by the European Research Council through the synergy grant UNIVERSE+, 101118787.
Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.